# Inductive-Deductive Strategy Reuse for Multi-Turn Instructional Dialogues

### v1

@@ 3.3 Model Training
We fine-tune LLaMA-2 Touvron et al. (2023) on $\mathcal{D}_{ins}$ to learn $P_{\phi}(\boldsymbol{U}|\mathbf{H},\boldsymbol{Q})$. In particular, the input is a prompt in Table 8 with the task description, the dialogue history $\mathbf{H}$, and the corresponding candidate $\boldsymbol{Q}$. The output is the selected instructional strategy and the instruction $\boldsymbol{U}$. We minimize the objective as follows:
@@
Benchmark Avg. # Turns Total # Dialogues
AlpacaEval 1 805
MT-Bench 2 80
MT-Bench++ 8 80
MT-Eval 6.96 168
Table 1: Statistics of the used benchmarks.
@@
We build a classification dataset $\mathcal{D}_{r}=\{(\mathbf{h}^{i},\boldsymbol{f}^{i},l^{i})_{i=1}^{R}\}$ based on $\mathcal{D}_{ins}$ and the corresponding $\mathcal{F}$, where $\mathbf{h}^{i}$ is a dialogue history and $\boldsymbol{f}^{i}$ is an instructional strategy. $l^{i}\in\{0,1\}$ denotes a matching label, which indicates that $\boldsymbol{f}^{i}$ is an appropriate strategy for $\mathbf{h}^{i}$ if $l^{i}=1$, otherwise $l^{i}=0$. For simplicity, the superscript $i$ will be omitted in the following. We fine-tune BERT Devlin et al. (2019) on $\mathcal{D}_{r}$ to learn $P_{\psi}(l|\mathbf{H},\boldsymbol{F})$ for correctly identifying the positive strategy from a set of negative strategies. We minimize the following objective,
@@
The positive is the golden strategy of the corresponding dialogue history, while the negatives are randomly selected from $\mathcal{F}$. One dialogue history has one positive and one negative strategy. Due to the BERT’s context length limitation, we only select the last-turn answer to represent the entire dialogue history. This ranker may have limited performance, but this lightweight model is sufficient for a coarse selection of candidates. Please see more analysis about this design in Appendix B.1.
@@ 4.1 Settings
For instruction generation, we use the human-machine instructional dialogues from ShareGPT to obtain instructional strategies, along with training the user simulator and the ranker in the induction stage. In the deduction stage, we randomly sample $10,000$ dialogues from unused ShareGPT and UltraChat datasets respectively, and then intercept the first-turn dialogues $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}\textquoteright_{0}\}$ as opening lines. Please see Appendix C for more details on data and method implementations.
@@ 4.2 Baselines
We compare IDEAS with a set of baselines: (1) Baize Xu et al. (2023b), which uses the instructions from opening lines as seeds and prompts GPT-4 to generate transcripts for both sides of the dialogues until a natural stopping point is reached. (2) UltraChat Ding et al. (2023), which adopts two separate GPT-4 APIs to generate dialogues based on the given opening lines. One plays the user’s role in generating instructions, and the other generates the answer. (3) ParrotChat Sun et al. (2023), which targets instructions extracted from human-machine dialogues as a learning goal and fine-tune a user simulator. The user simulator and GPT-4 iteratively interact to generate instructional dialogues. (4) SkillGen, which takes a similar approach to training the user simulator. The difference is that the user simulator is fine-tuned to sequentially generate a high-level instructional strategy and instruction for the given dialogue history. For a fair comparison, all experimental setups of baselines are consistent with IDEAS. Those baselines using user simulators also introduce our reflection module for quality control. In addition, we compare the above models with GPT-4 and ChatGPT as chat models to clarify the room for improvement.
@@
Appr. Coh. Depth Insight Diversity Appr. Coh. Depth Insight Diversity
Baize ${}^{\ \,}$ ${}^{\ \,}$ 2.65 2.49 2.11 ${}^{\ \,}$ ${}^{\ \,}$ 2.77 2.59 2.28
UltraChat ${}^{\ \,}$ ${}^{\ \,}$ 3.18 3.04 3.05 ${}^{\ \,}$ ${}^{\ \,}$ 3.42 3.19 3.17
ParrotChat 7.89 7.71 5.19 4.62 4.46 7.74 7.47 5.89 5.29 5.16
@@
IDEAS ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$
ShareGPT ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$
Table 2: Automatic evaluation of the generated instructions by different methods on different dialogue scenarios, including ShareGPT and UltraChat. The bottom row corresponds to the human-machine instructional dialogues in $\mathcal{D}_{ins}$. Significance tests between IDEAS and baselines are performed using t-test. $\ddagger$ indicates $p$-value < 0.01.
@@ 4.3 Evaluation Metrics
Data Evaluation.
We follow Ding et al. (2023) to design the following metrics to evaluate generated instructions. These metrics are implemented by prompting gpt-4-1106-preview to output scores, which can be found in Table 10. (1) Appropriateness (Appr.), which means whether the instruction is appropriate with the given dialogue history and should not be detached from the history or topic. (2) Coherence (Coh.), which means whether the instruction is logically coherent to the above instructions or the answers in the dialogue history. (3) Depth, which means whether the instruction expands the topic or explores more details in the dialogue history. (4) Insight, which means whether the instruction can bring new understanding, stimulate thought, improve the dialogue’s quality, or help unearth more valuable information, instead of repeating known content. (5) Diversity, which means whether the instruction is different from the above instructions. The difference is reflected in the instruction type. The rating scale is of $1$ to $10$, in which $1$ means worst and $10$ best.
@@
We use the following widely-used benchmarks to evaluate chat models. (1) AplacaEval (AE) Dubois et al. (2023), which is an automatic evaluator for instruction-following language models. (2) MT-Bench (MB) Zheng et al. (2023), which contains two-turn instructional dialogues that evaluate a chatbot’s multi-turn conversational and instruction-following ability; (3) MT-Bench++  (MB++) Sun et al. (2023), which expands the dialogues in MT-Bench to create an eight-turn evaluation dataset. (4) MT-Eval (ME) Kwan et al. (2024), which is a comprehensive benchmark to evaluate multi-turn dialogue abilities. The statistics of these benchmarks are shown in Table 1.
@@ 5 Results and Discussion
5.1 Data Evaluation
We first evaluate the generated instructions. To investigate the consistency between GPT-4 and human annotators, we randomly select $500$ instructions for human evaluation. Specifically, We employ three annotators to rate the instructions and present the same evaluation instruction as GPT-4. The consistency is measured via the Fleiss’s kappa $\kappa$ Randolph (2005). The $\kappa$ values for Appropriateness, Coherence, Depth, Insight, Diversity are 0.71, 0.67, 0.59, 0.53, and 0.64 respectively.
The results are shown in Table 2, indicating that our generated instructions outperform all the baselines on the key optimized metrics. We further observe that: (1) Our generated instructions achieve scores similar to the human-machine dialogues on all metrics, indicating that our generated instructions are high-quality. We also present the constructed dialogues of ParrotChat and IDEAS in Table 11 and 12. (2) IDEAS achieves higher scores on all metrics than ParrotChat, indicating the effectiveness of instructional strategy reuse via induction and deduction for instruction generation. (3) SkillGen achieves better scores of Diversity, Depth, and Insight than ParrotChat, which indicates that the explicit modeling of the complex dialogue flow is effective for posing more diverse, in-depth, and insightful instructions. (4) Baize achieves the highest scores on Appropriateness and Coherence. We speculate that GPT-4 manages the generation of the entire dialogue, which may perform better in generating appropriate and coherent instructions. Overall, IDEAS achieves relatively good scores in terms of Appropriateness and Coherence compared to other methods using user simulators.
@@ 5.1 Data Evaluation
Baize 58.39 6.15 5.89 7.02
UltraChat ${}^{\ \,}$ 6.78 6.83 7.01
ParrotChat 77.89 6.71 6.78 7.07
@@
Baize 44.58 6.06 6.28 6.43
UltraChat ${}^{\ \,}$ 6.30 6.67 6.40
ParrotChat 74.66 6.38 6.57 6.54
@@ 5.2 Model Evaluation
We further evaluate the benefit of the constructed dialogues on the downstream chat models. The results are shown in Table 3, which indicates that IDEAS outperforms all the baselines on almost all the benchmarks. This confirms the effectiveness of generating diverse, in-depth, and insightful instructions. We further observe that: (1) IDEAS achieves higher scores for all multi-turn benchmarks compared to baselines, especially ParrotChat. This demonstrates that inductive-deductive strategy reuse is effective for improving multi-turn dialogue and instruction-following abilities. (2) UltraChat achieves the highest score on AlpacaEval which focuses on single-turn instruction-following ability, but performs relatively poorly on the multi-turn benchmarks. This indicates that the dialogues generated by two role-played LLMs are more like the concatenation of single-turn data, i.e., insufficient depth of multi-turn dialogues.
@@
5.3 Further Discussion
We also investigate the effect of each component of IDEAS, the impact of the abstraction level, and the impact of the amount of instructional dialogues.
@@ 5.3 Further Discussion
We perform the following ablation tests to validate the effect of each component: (1) Do not guarantee the quality of each instruction via quality control (w/o Reflection). (2) Directly generate a strategy without the candidate for selection (w/o CandGen). (3) Randomly choose $W$ strategies from $\mathcal{F}$ as the candidate without ranking (w/o Ranker). (4) Providing the top-1 strategy via ranker to the user simulator without secondary selection (w/o CandTop1). (5) Providing a strategy that randomly selects from the candidate without secondary selection (w/o CandRand). (6) Selecting $W$ original strategies from $\mathcal{F}_{o}$ to replace the high-level strategies as the candidate (w/o Abstraction). The results are shown in Table 4. We observe that ablating each component brings varying degrees of performance drop. This demonstrates the necessity of designing all these components.
Threshold MT-Bench MT-Bench++ MT-Eval
$\epsilon=0.4$ 6.76 ${}^{\ \,}$ 7.18
$\epsilon=0.5$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$
$\epsilon=0.6$ 6.70 6.82 7.16
Table 5: Performance changes on the chat model under the ShareGPT dialogue scenario by providing different $\epsilon$, i.e., different abstraction levels. $\dagger$ and $\ddagger$ indicate $p$-value < 0.05 and 0.01 (significance tests via t-test).
The Impact of Abstraction Level $\epsilon$.
We set the similarity threshold $\epsilon$ to different values, i.e., $0.4$, $0.5$, $0.6$, for assessing the impact of the abstraction level of instructional strategies. The results are shown in Table 5. We can observe that the scores of the chat model with the threshold $0.5$ reach a peak on all benchmarks, and drop as the level of abstraction increases (i.e., $0.4$). This suggests that moderate abstraction that removes irrelevant details and distills high-level strategies is optimal. We speculate that moderate abstraction can guide the user simulator to focus on key patterns in those dialogue histories that can use the same high-level strategy. This helps the user simulator recognize key patterns in a newly given dialogue history to select an appropriate strategy without being distracted by irrelevant details. However, excessive abstraction may reduce the accuracy of generalization and make it more difficult for the user simulator to focus on key common patterns.
Figure 3: Performance changes on chat models respectively by providing different amounts of instructional dialogues generated by IDEAS and ParrotChat.
@@
We select $0.5$x, $1$x, and $1.5$x the amount of generated dialogues to assess the impact of providing more dialogues and compare IDEAS with the baseline, i.e., ParrotChat. Note that $1$x represents that 1*10K generated dialogues are selected. Therefore, we select an additional 5K opening lines from unused ShareGPT data and then apply IDEAS and ParrotChat to generate more dialogues. The results are shown in Figure 3. We can observe that: (1) The scores of ParrotChat on all benchmarks reach a peak at 1x and no increase afterward. We speculate that the implicit modeling of complex dialogue flows makes ParrotChat tend to generate general instructions, resulting in limited diversity. Similar dialogues at high amounts would not positively affect training. (2) However, the scores of IDEAS keep increasing from 0.5x to 1.5x. This indicates that IDEAS can generate diverse dialogues, further improving the performance of the downstream chat model.

### v2

@@ 3.3 Model Implementation
We follow Kong et al. (2023b) and Sun et al. (2023) to choose the fine-tuned open-source LLM as the user simulator due to its lightweight and free. Specifically, we fine-tune LLaMA-2 Touvron et al. (2023) on human-machine instructional dialogues $\mathcal{D}_{ins}$ to learn $P_{\phi}(\boldsymbol{U}|\mathbf{H},\boldsymbol{Q})$. The input is a prompt in Table 11 with the task description, the dialogue history $\mathbf{H}$, and the corresponding candidate $\boldsymbol{Q}$. The output is the selected instruction strategy and the instruction $\boldsymbol{U}$. We minimize the objective as follows:
@@
We build a classification dataset $\mathcal{D}_{r}=\{(\mathbf{h}^{i},\boldsymbol{f}^{i},l^{i})_{i=1}^{R}\}$ based on $\mathcal{D}_{ins}$ and the corresponding $\mathcal{F}$, where $\mathbf{h}^{i}$ is a dialogue history and $\boldsymbol{f}^{i}$ is an instruction strategy. $l^{i}\in\{0,1\}$ denotes a matching label, which indicates that $\boldsymbol{f}^{i}$ is an appropriate strategy for $\mathbf{h}^{i}$ if $l^{i}=1$, otherwise $l^{i}=0$. For simplicity, the superscript $i$ will be omitted in the following. We fine-tune BERT Devlin et al. (2019) on $\mathcal{D}_{r}$ to learn $P_{\psi}(l|\mathbf{H},\boldsymbol{F})$ for correctly identifying the positive strategy from a set of negative strategies. We minimize the following objective,
@@
For a given dialogue history $\mathbf{h}^{i}$ in $\mathcal{D}_{ins}$, the positive is the high-level strategy $\boldsymbol{f}^{i}$ obtained from the subsequent instruction of $\mathbf{h}^{i}$ in $\mathcal{D}_{ins}$, while the negative is randomly selected from $\mathcal{F}$. One dialogue history has one positive strategy and one negative strategy. Due to the BERT’s context length limitation, we only select the last-turn answer to represent the entire dialogue history. This ranker may have limited performance, but this lightweight model is sufficient for a coarse selection of candidates. Please see more analysis about this design in Appendix A.1.
@@ 4.1 Settings
The basic data used for all experiments are collected by Sun et al. (2023). For instruction generation, we use the human-machine instructional dialogues from ShareGPT to obtain instruction strategies, along with training the user simulator and the ranker in the induction stage. In the deduction stage, we randomly sample 10K dialogues from unused ShareGPT and UltraChat Ding et al. (2023) datasets, and then intercept the first-turn $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}\textquoteright_{0}\}$ as opening lines. Please see Appendix B for more details on data and method implementations.
@@ 4.2 Baselines
We compare IDEAS with a set of baselines: (1) Self-Chat Xu et al. (2023b), which uses the instructions from opening lines as seeds and prompts GPT-4 to generate transcripts for both sides of the dialogues until a natural stopping point is reached. (2) Iterative Self-Chat (Iter Self-Chat) Ding et al. (2023), which adopts two separate GPT-4 to generate dialogues based on the given opening lines. One simulates the use in generating instructions, and the other generates the answer. (3) Parrot-Ask Sun et al. (2023), which fine-tunes LLaMA-2 as the user simulator. It takes the dialogue history as the input to generate instructions. (4) SkillGen, which is a variant of IDEAS. The input is the given dialogue history and the output is the sequentially generated instruction strategy and instruction. The difference is that IDEAS provides strategy candidates for the user simulator to select rather than generate the appropriate strategy. Note that the system agents in all baselines adopt GPT-4. Those baselines using fine-tuned LLaMA-2 as user simulators also introduce our reflection module for quality control. For a fair comparison, all experimental setups of baselines are consistent with IDEAS. We also test three different abstraction level in Eq. 4, i.e., $\epsilon=0.4,0.5,0.6$. In addition, we supplement GPT-4 and ChatGPT as chat models to clarify the room for improvement.
@@
Appr. Coh. Dep. Ins. Div. Appr. Coh. Dep. Ins. Div.
Self-Chat ${}^{\ \,}$ ${}^{\ \,}$ 2.65 2.49 2.11 ${}^{\ \,}$ ${}^{\ \,}$ 2.77 2.59 2.28
Iterative Self-Chat ${}^{\ \,}$ ${}^{\ \,}$ 3.18 3.04 3.05 ${}^{\ \,}$ ${}^{\ \,}$ 3.42 3.19 3.17
Parrot-Ask 7.89 7.71 5.19 4.62 4.46 7.74 7.47 5.89 5.29 5.16
@@
IDEAS $\epsilon=0.6$ 9.12 8.91 6.57 5.99 5.51 9.30 9.08 7.41 6.77 6.68
IDEAS $\epsilon=0.5$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$
IDEAS $\epsilon=0.4$ 9.07 8.86 6.59 5.92 5.59 9.26 8.96 7.38 6.63 6.56
$\mathcal{D}_{ins}$ (human) ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$ ${}^{\ \,}$
Table 1: Automatic evaluation of the generated instructions by different methods on ShareGPT and UltraChat dialogue scenarios. The bottom row corresponds to the human-machine instructional dialogues in $\mathcal{D}_{ins}$. Significance tests between IDEAS ($\epsilon=0.5$) and baselines are performed using t-test. $\ddagger$ indicates $p$-value < 0.01.
@@ 4.3 Evaluation Metrics
Instruction Evaluation.
Efficient instruction evaluation about depth and diversity is significantly understudied. Thus, we follow Ding et al. (2023) to design five metrics to evaluate generated instructions and implement each based on “LLMs as Judges” Zheng et al. (2024b) by prompting the top-tier LLM, i.e., gpt-4-1106-preview to output scores  [^3] GPT-4 prefers to generate instructions for itself as the system agent, which may lead to inflated scores for Self-Chat and iter Self-Chat. However, it does not affect the ranking of these methods in generating diverse and in-depth instructions from the experimental results in Section 5.1 since Self-Chat and Iter Self-Chat achieve the lowest scores on related metrics.. The prompt is shown in Table 13. (1) Appropriateness (Appr.), which means whether the instruction is appropriate with the given dialogue history and should not be detached from the history. (2) Coherence (Coh.), which means whether the instruction is logically coherent to the given dialogue history. (3) Depth (Dep.), which means whether the instruction expands the topic or explores more details in the dialogue history. (4) Insight (Ins.), which means whether the instruction can bring new understanding, stimulate thought, lead to deeper interaction, or help unearth more valuable information, instead of repeating known content. (5) Diversity (Div.), which means whether the instruction is different from the previous instructions. The difference is reflected in the instruction type. The rating scale is of $1$ to $10$, in which $1$ means worst and $10$ best.
@@
We use the following widely-used benchmarks to evaluate chat models. (1) AlpacaEval (AE) Dubois et al. (2023), which is an automatic evaluator for the single-turn instruction-following ability. (2) MT-Bench (MB) Zheng et al. (2023), which contains two-turn instructions that evaluate a chatbot’s multi-turn conversational and instruction-following ability; (3) MT-Bench++  (MB++) Sun et al. (2023), which expands the dialogues in MT-Bench to create an eight-turn evaluation dataset. (4) MT-Eval (ME) Kwan et al. (2024), which is a comprehensive benchmark to evaluate multi-turn dialogue abilities. The statistics of these benchmarks are shown in Table 5.
@@ 5 Results and Discussion
5.1 Instruction Evaluation
We first evaluate the generated instructions. To investigate the consistency between GPT-4 Judger and human annotators, we randomly select $500$ instructions for human evaluation. Specifically, We employ three annotators to rate the instructions and present the same evaluation instruction as GPT-4. The consistency is measured via the Fleiss’s kappa $\kappa$ Randolph (2005). The $\kappa$ values for Appropriateness, Coherence, Depth, Insight, Diversity are 0.71, 0.67, 0.59, 0.53, and 0.64 respectively.
The results are shown in Table 1, indicating that our generated instructions outperform all the baselines on the key optimized metrics. We further observe that: (1) Our generated instructions achieve scores similar to $\mathcal{D}_{ins}$ on all metrics, indicating that our generated instructions are high-quality. We present the constructed dialogues in Table 14 and 15. (2) IDEAS achieves higher scores of Dep., Ins. and Div., indicating that IDEAS can generate more diverse and in-depth instructions. In particular, Parrot-Ask vs. IDEAS shows the effectiveness of inductive-deductive strategy reuse. (3) SkillGen vs. IDEAS indicates the effectiveness of the provided candidate in strategy utilization. (4) Self-Chat achieves the highest scores on Appr. and Coh.. We speculate that GPT-4 manages the generation of the entire dialogue, which may perform better in generating appropriate and coherent instructions. Besides, the scores are close to the full score. This is also partly because GPT-4 tends to give higher scores than the actual scores when scoring the text that it generates. Overall, IDEAS achieves relatively good scores of Appr. and Coh. (5) The threshold $\epsilon=0.5$ achieves higher scores on all the metrics. Please refer to Section 5.3 for further analysis. Thus, $0.5$ can be used for building high-quality instructional dialogues.
Avg.#Turns Avg.#Tokens
Self-Chat 3.10 28.53
Iterative Self-Chat 8.08 31.38
Parrot-Ask 9.32 36.85
SkillGen 9.37 36.29
IDEAS 9.33 36.90
Table 2: Data statistics. Avg.#Turns and Avg.#Tokens denote the average turns of constructed instructional dialogues and the average length of generated instructions.
@@ 5.1 Instruction Evaluation
Self-Chat 58.39 6.15 5.89 7.02
Iter Self-Chat ${}^{\ \,}$ 6.78 6.83 7.01
Parrot-Ask 77.89 6.71 6.78 7.07
@@
Self-Chat 44.58 6.06 6.28 6.43
Iter Self-Chat ${}^{\ \,}$ 6.30 6.67 6.40
Parrot-Ask 74.66 6.38 6.57 6.54
@@ 5.2 Model Evaluation
We further evaluate the benefit of the constructed dialogues on the downstream chat models. Table 2 first shows data statistics. The results are shown in Table 3, which indicates that IDEAS outperforms all the baselines on almost all the benchmarks. This confirms the effectiveness of generating diverse and in-depth instructions. We further observe that: (1) IDEAS achieves higher scores for all multi-turn benchmarks compared to baselines, especially Parrot-Ask. This demonstrates that strategy reuse is effective for improving multi-turn dialogue and instruction-following abilities. (2) Iter Self-Chat achieves the highest score on AlpacaEval which focuses on the single-turn instruction-following ability, but performs relatively poorly on the multi-turn benchmarks. This indicates that the instructions generated by the role-played GPT-4 in multi-turn interaction are weakly correlated, i.e., insufficient depth of multi-turn instructions.
5.3 Further Discussion
We investigate the impact of the abstraction level, the effect of each component of IDEAS, and the impact of the amount of instructional dialogues.
The Impact of Abstraction Level.
We set the similarity threshold $\epsilon$ to different values, i.e., $0.4$, $0.5$, $0.6$, for assessing the impact of the abstraction level of instruction strategies. The results are shown in Table 1. We can observe that the scores of the generated instructions with the threshold $0.5$ reach a peak on all metrics, and drop as the level of abstraction increases (i.e., $0.4$). This suggests that moderate abstraction that removes unnecessary details and distills high-level strategies is optimal. We speculate that insufficient abstraction leads to strategies that involve details related to specific dialogues, i.e., are overly specific. However, these details may lead to strategies that do not fully match a new dialogue history. Thus, these seemingly appropriate strategies are difficult to effectively utilize by user simulators, which tend to generate more general instructions. Furthermore, excessive abstraction makes one strategy applicable to more dialogue histories, i.e., one strategy can guide multiple dialogue flows. This still results in diminished diversity and insufficient depth.
@@ 5.3 Further Discussion
We perform the following ablation tests to validate the effect of each component: (1) Do not guarantee the quality of each instruction via quality control (w/o Reflection). (2) Directly generate a strategy without the candidate for selection (w/o CandGen). (3) Randomly choose $W$ strategies from $\mathcal{F}$ as the candidate without ranking (w/o Ranker). (4) Provide the top-1 strategy via ranker to the user simulator without secondary selection (w/o CandTop1). (5) Provide a strategy that randomly selects from the candidate without secondary selection (w/o CandRand). (6) Select $W$ original strategies from $\mathcal{F}_{o}$ to replace the high-level strategies as the candidate (w/o Abstraction). The results are shown in Table 4. We observe that ablating each component brings varying degrees of performance drop. This demonstrates the necessity of designing all these components.
@@
We select $0.5$x, $1$x, and $1.5$x the amount of generated dialogues to assess the impact of providing more multi-turn instructional dialogues and compare IDEAS with Parrot-Ask. Note that $1$x represents that 1*10K generated multi-turn instructional dialogues are selected. Therefore, we select an additional 5K opening lines from unused ShareGPT data Sun et al. (2023) and then apply IDEAS and Parrot-Ask to generate more dialogues. The results are shown in Figure 3. We can observe that: (1) The scores of Parrot-Ask on all benchmarks reach a peak at 1x and no increase afterward. We speculate that the Parrot-Ask struggles to implicitly capture complex rules, which leads to more generic instructions. Further, similar dialogues at high amounts would not positively affect training Ou et al. (2022). (2) In contrast, the scores of IDEAS keep increasing from 0.5x to 1.5x. This indicates that the multi-turn instructional dialogue data constructed by IDEAS is more diverse, further improving the performance of the downstream chat model.
Figure 3: Performance changes on chat models respectively by providing different amounts of instructional dialogues generated by IDEAS and Parrot-Ask.